OpenStudy (anonymous):

The polynomial of degree 3, P(x), has a root of multiplicity 2 at x=4 and a root of multiplicity 1 at x=−5. The y-intercept is y=−32. Find a formula for P(x).

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@wio

jimthompson5910 (jim_thompson5910):

x = 4, x = 4, or x = -5 x - 4 = 0 or x - 4 = 0 or x + 5 = 0 k(x-4)(x-4)(x+5) = 0 k(x-4)(x^2 + x - 20) = 0 k[ x(x^2 + x - 20) - 4(x^2 + x - 20) ] = 0 k( x^3 + x^2 - 20x - 4x^2 - 4x + 80 ) = 0 k( x^3 - 3x^2 - 24x + 80 ) = 0 y = k( x^3 - 3x^2 - 24x + 80 ) Now plug in x = 0 and y = -32 then solve for k

OpenStudy (anonymous):

ok thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

i got k=-5/2 and then put it in and got it wrong

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = k( x^3 - 3x^2 - 24x + 80 ) -32 = k( (0)^3 - 3(0)^2 - 24(0) + 80 ) -32 = 80k 80k = -32 k = -32/80 k = -2/5 so you flipped it somehow

OpenStudy (anonymous):

my bad